The following program is used to write data to a printer using the SPP protocol showed below.

void PutChar(unsigned char c)
{
while ((PTJ & 0x02) !=0 );

PTH = c;

PTJ = PTJ & ~0x01;
NOP;
//....
NOP;
PTJ = PTJ | 0x01;
}


DB-Do STRB# ACK# BUSY 8 min, min. 5 min.min. »

How many NOP statements must be used to insure the STRB# line is low for the minimum amount of time if n = 0.5 ?s and the processor takes 5clock tick(s) per instruction with a clock of 10.0 MHz?
Answer #1

We can see that we have to make it low for 0.5 sec and each clock cycle is 1/10*10^6=100 ns

where ns is nano secs

Then each instruction takes 5*100=500 ns

And Hence 0.5=500ns*n where n is the number of nop instructions

Hence we have 0.5/5 10 =10^6 instructions

Or 10^6 NOP statements